OpenStudy (anonymous):

The probability of an event occurring is 23/50. What are the odds that the event will occur?

OpenStudy (anonymous):

if the probability in favor is p, then the odds in favor are \[\frac{p}{1-p}\] i.e. probabilty in favor divided by the probability against. in this case you get \[\frac{23}{27}\]

OpenStudy (anonymous):

think of it easy as 23 out of 50 chances to win and therefore 27 out of 50 to lose.

OpenStudy (amistre64):

the odds are 46%

OpenStudy (amistre64):

23/50 that it occurs. the odds that it WILL occur are 46% right?

OpenStudy (anonymous):

really? in that case i made a mistake, but i don't think so.

OpenStudy (anonymous):

oh no that is probability, not odds.

OpenStudy (amistre64):

ahhh... like saying 5 to 2 or something right?

OpenStudy (anonymous):

if the probability i win is \frac{2}{3}\] then the odds i win are two to one.

OpenStudy (anonymous):

\[\frac{p}{1-p}\]

OpenStudy (amistre64):

chance occur to chance not occur is an odd then right?

OpenStudy (anonymous):

right like saying 5 to 2. exactly

OpenStudy (anonymous):

yes right.

OpenStudy (anonymous):

not really used much in math, but sometimes. gamblers' ruin problem. used mostly at the track.